### v1

# QLLM: Do We Really Need a Mixing Network for Credit Assignment in Multi-Agent Reinforcement Learning?

Abstract
Credit assignment has remained a fundamental challenge in multi-agent reinforcement learning (MARL). Previous studies have primarily addressed this issue through value decomposition methods under the centralized training with decentralized execution paradigm, where neural networks are utilized to approximate the nonlinear relationship between individual Q-values and the global Q-value. Although these approaches have achieved considerable success in various benchmark tasks, they still suffer from several limitations, including imprecise attribution of contributions, limited interpretability, and poor scalability in high-dimensional state spaces. To address these challenges, we propose a novel algorithm, QLLM, which facilitates the automatic construction of credit assignment functions using large language models (LLMs). Specifically, the concept of TFCAF is introduced, wherein the credit allocation process is represented as a direct and expressive nonlinear functional formulation. A custom-designed coder-evaluator framework is further employed to guide the generation, verification, and refinement of executable code by LLMs, significantly mitigating issues such as hallucination and shallow reasoning during inference. Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, QLLM exhibits strong generalization capability and maintains compatibility with a wide range of MARL algorithms that utilize mixing networks, positioning it as a promising and versatile solution for complex multi-agent scenarios.
123
1 Introduction
Recent advances in cooperative multi-agent reinforcement learning (MARL) [23, 3] have led to widespread applications in domains such as assembly lines [13], logistics distribution [42], autonomous driving [49], and swarm robotics [24]. In such scenarios, multiple agents learn to interact with the environment collaboratively to accomplish designated cooperative tasks. Given that agents receive a shared team reward, it is essential to accurately assess the individual contribution of each agent to facilitate effective training. A key challenge in this context is the credit assignment problem, where inaccurate attribution of individual contributions may result in the emergence of so-called "lazy agents" [18].
To address the credit assignment problem in multi-agent systems, a variety of solutions have been proposed in the academic community. Attention-based credit assignment methods [33, 22] incorporate a learnable mechanism that adaptively allocates credit by computing attention weights over agents or their interactions. By explicitly modeling inter-agent dependencies, these methods enable the network to selectively attend to the most influential agents or environmental factors when estimating individual contributions. This dynamic weighting scheme enhances the expressiveness of the credit assignment process, particularly in heterogeneous or partially observable environments.
Policy gradient approaches [1] tackle credit assignment from a differentiable optimization perspective. Each agent’s policy is updated based on gradients computed from shared rewards, typically combined with baselines or variance reduction techniques to improve learning stability. These methods can naturally handle continuous action spaces and stochastic policies, where credit assignment is implicitly achieved by estimating each agent’s impact on the long-term return, either through joint trajectory sampling or individualized advantage estimation. In addition, value decomposition methods [35, 38] have become increasingly popular. These methods address the credit assignment problem by constructing a joint value function that can be decomposed into individual utility functions associated with each agent. The decomposition is designed to ensure that the optimal joint policy can be recovered from the individual components, allowing for centralized training while preserving decentralized execution. These methods rely on specific structural assumptions that maintain the consistency between the global value and its local components, thereby enabling efficient coordination and scalable training.
However, most existing approaches predominantly rely on neural networks, either explicitly or implicitly, to perform credit assignment for individual agents. A representative example of this is the mixing network employed in value decomposition methods. These approaches often suffer from high training costs, limited generalization ability, and a lack of sematic interpretability [17, 29]. With the advancement of large language models (LLMs) [12], new paradigms have emerged to address many long-standing challenges in MARL. By leveraging extensive pre-trained knowledge [25], superior generalization capacity[2], and powerful code generation abilities [39], LLMs have demonstrated potential as versatile auxiliary components in reinforcement learning pipelines. Recent studies have shown the effectiveness of LLMs in generating dense, agent-specific rewards based on natural language descriptions of tasks and overall team goals. Empirical results reveal that such LLM-driven reward shaping leads to faster convergence and improved policy performance over existing baselines. Furthermore, the generalization and multitask learning capabilities of LLMs allow agents to quickly adapt to novel tasks with limited training samples, thus improving sample efficiency [30]. In tasks involving human interaction, LLMs can generate responses that more closely align with human expectations, enhancing interaction quality [48]. Given the extensive applications of LLMs in the field of reinforcement learning, we propose to leverage them to address the credit assignment problem in multi-agent systems.
In this paper, we introduce QLLM, a novel value decomposition method that eliminates the need for a mixing network. To support its construction, we first propose a general coder-evaluator framework for LLM-based code generation. Within this framework, LLMs act as both a code generator and a code evaluator [5]. Upon receiving task and code generation instructions, the code generator produces a non-linear function called Training-Free Credit Assignment Function (TFCAF), which serves as a direct substitute for the mixing network in algorithms such as QMIX, VMIX, and RIIT [28, 34, 10]. Generated directly by the LLMs, this credit assignment function offers several advantages, including no need for training, immunity to high-dimensional state spaces, and enhanced semantic interpretability. Additionally, a code verification mechanism has been designed to ensure that the generated code runs correctly while mitigating hallucination issues [21] commonly encountered by LLMs. This framework significantly reduces the probability of errors when LLMs process complex textual prompts and enables the generation of a reliable TFCAF without direct interaction with the environment, thereby allowing it to replace the mixing network. Since LLMs modify only the mixing network in MARL algorithms, the TFCAF is applicable to almost any MARL algorithm that relies on mixing networks.
The main contributions of this study are summarized as follows:
1.
We propose the coder-evaluator framework for LLM-based code generation, which enables LLMs to generate a reliable credit assignment function under zero-shot prompting and without any human feedback, referred to as TFCAF. This framework is generalizable and can be extended to other reinforcement learning scenarios enhanced by LLMs.
2.
QLLM is proposed by combining TFCAF and value function decomposition while remaining compatible with most MARL algorithms that employ mixing networks for credit assignment. Unlike existing methods, QLLM can accurately infer each agent’s contribution without interacting with the environment.
3.
Extensive experiments demonstrate that QLLM outperforms existing credit assignment algorithms, particularly in scenarios involving high-dimensional state spaces. Moreover, the integration of LLMs endows the contribution assignment among agents with semantic interpretability.
2 Related Work
2.1 Credit Assignment in MARL
In recent years, several algorithms have been proposed to address the credit assignment problem. And policy-based approaches [45, 47] were among the earliest introduced. MADDPG [19] extends the DDPG [15] framework to multi-agent settings by incorporating centralized training with decentralized execution. This approach allows each agent to learn policies that account for the actions of other agents during training, thereby addressing the non-stationarity inherent in multi-agent environments. MADDPG utilizes an actor-critic architecture with centralized critics, where the centralized critic has access to global information, enabling more effective learning in mixed cooperative-competitive scenarios. Similarly, COMA [6] leverages a counterfactual baseline to compute each agent’s advantage function, following the same actor-critic framework as MADDPG. MAAC [11] further enhances policy-based methods by introducing an attention mechanism into the critic network, leading to more precise credit assignment. However, most of these algorithms are not directly applicable to tasks with shared rewards.
Another widely adopted paradigm is the value decomposition approach [28, 40, 27, 41], which explicitly decomposes the globally shared return. As one of the earliest methods in this category, VDN [35] models the global value function as a linear summation of individual value functions. While simple and computationally efficient, VDN struggles to capture complex cooperative relationships due to its linear decomposition. QMIX [28] improves upon VDN by using a monotonic mixing network to non-linearly combine individual value functions, allowing for more expressive representations. However, the assumption of monotonicity between individual and global values may restrict its flexibility in highly complex environments. QTRAN [32] and QPLEX [37] mitigate this limitation by relaxing the monotonicity constraint, enabling the model to learn richer representations, albeit at the cost of increased computational complexity. LICA [50] adopts an alternative approach by introducing an independent credit assignment network that directly derives individual contributions from global information. Unlike methods focused solely on value decomposition, LICA explicitly models individual contributions, leading to more accurate credit assignment in cooperative tasks. RIIT [10] goes further by modeling implicit influence relationships between agents, enabling better credit assignment in tasks with long-term dependencies. Nevertheless, both RIIT and LICA continue to employ a mixing network structure analogous to that used in QMIX, in order to establish the mapping from individual agent Q-values to the global Q-value. Although the mixing network is a critical component for implementing credit assignment, it suffers from limited interpretability and slow convergence during training. Despite the advantages of value decomposition methods, achieving efficient and precise credit assignment in multi-agent systems remains a challenging open research problem.
2.2 LLM-enhanced RL
Leveraging their large-scale pretraining and strong generalization capabilities, LLMs have been increasingly integrated into Reinforcement Learning to enhance performance in multitask learning, sample efficiency, and high-level task planning [46, 14]. In the SMAC-R1 framework [4], agents utilize LLMs to generate decision tree code by providing task descriptions. The model further improves the quality and interpretability of the decision tree through self-reflection, guided by feedback from environment-provided rewards. Additionally, LLMs can serve as generators to create high-fidelity multimodal world models, reducing the cost of learning in real-world environments. They can also generate natural language rationales to explain agent behaviors, improving the transparency of decision-making processes. For instance, the Reflexion framework [31] introduces a reinforcement mechanism guided by natural language feedback, without requiring updates to the model parameters. Agents reflect on task-related feedback signals through language and store them in contextual memory, enabling improved decision-making in future interactions. Moreover, LLMs can explicitly or implicitly shape reward signals for RL tasks. The EUREKA framework [20] achieves superhuman-level reward function coding through evolutionary search and policy feedback, empowering agents to execute more complex behaviors.
These examples demonstrate the powerful information processing and code generation capabilities of LLMs. Therefore, some researchers have leveraged LLMs to tackle the challenges of temporal and agent-level credit assignment in MARL. Qu et al. [26] were the first to introduce LLMs into the credit assignment problem. They proposed the notion of latent rewards, making LLMs to capture subtle cooperative relationships among agents and to decompose global rewards accordingly. This method effectively addresses the challenge of multi-agent credit assignment under sparse rewards and outperforms existing baseline algorithms. However, it suffers from instability in code generation and may produce latent rewards that contain redundant or difficult-to-interpret components. Additionally, Lin et al. [16] employed LLMs to learn potential-based reward functions across multiple queries, allowing for a more accurate estimation of individual agent contributions. These methods primarily focus on decomposing the global reward across agents but remain limited to settings compatible with single-agent reinforcement learning, which constrains their scalability and performance. Therefore, we propose QLLM, a multi-agent credit assignment algorithm based on the coder-evaluator framework, to address the aforementioned limitations.
3 Preliminaries
3.1 Decentralized Partially Observable Markov Decision Process
In this work, we focus on a fully cooperative MARL setting that can be formalized as a decentralized partially observable Markov decision process (Dec-POMDP). A Dec-POMDP is typically defined by a tuple . Here, agents indexed by operate in an environment with true underlying state . At each timestep, agent selects an action based solely on its private observation , which is sampled according to the observation function . The joint action of all agents is denoted by .
The environment evolves according to the transition probability function , which specifies the likelihood of transitioning to state given the current state and joint action . All agents cooperate to maximize a shared reward signal defined by the function . The objective in this setting is to learn joint policies that maximize the expected cumulative discounted reward , where is a discount factor determining the weight of future rewards.
A key challenge in Dec-POMDPs lies in the credit assignment problem. Since all agents share the same reward, it is difficult to attribute individual contributions. Moreover, as in other multi-agent settings, the environment is non-stationary from each agent’s perspective due to the concurrent learning processes of other agents, posing significant challenges to coordination and decentralized control.
3.2 Value Decomposition Methods
Traditional value decomposition-based algorithms employ a mixing network to combine the local Q-value functions of individual agents into a global Q-value function . The mixing network serves as a differentiable function approximator that aggregates individual agent utilities while ensuring that the global Q-value adheres to the Individual-Global-Maximum (IGM) constraint. The IGM principle [9] requires that the joint action that maximizes the global action-value function must correspond to the combination of each agent’s individually optimal action. Formally, the IGM principle can be expressed as:
where represents the joint action-observation histories of all agents. This equation implies that the maximization of the global Q-value function is aligned with the maximization of each agent’s local Q-value function, enabling decentralized execution while preserving global optimality. Building upon this, Qatten [43] provides a theoretical extension, showing that can be expressed as a linear combination of the individual functions:
denotes a coefficient that depends on the state , where is the number of attention heads. Theoretically, these coefficients can be derived from a higher-order functional expansion of with respect to , where the contribution of each term diminishes rapidly with increasing derivative order. This indicates that the influence of higher-order interaction terms on is sufficiently small to be ignored or effectively absorbed into a bias term , allowing for a simplified model formulation and reduced computational complexity. Meanwhile, the theoretical insight in Equation (2) lays the groundwork for the subsequent development of TFCAF.
4 Method
This section presents a detailed discussion of the implementation of each component of QLLM. We first introduce the concept of TFCAF, a training-free credit assignment function capable of accurately assigning credit to individual agents. Leveraging the generative capabilities of LLMs, TFCAF is constructed through prompt-based interaction with LLMs. However, when applied to complex tasks, LLMs typically encounter two fundamental challenges, hallucination [36] and a lack of rigorous reasoning. These limitations motivate the design of our proposed coder-evaluator framework. By integrating this framework, QLLM is able to generate highly accurate TFCAF functions for credit assignment in multi-agent settings, offering an effective alternative to the traditional mixing network used in value decomposition approaches.
4.1 Training-Free Credit Assignment Function
As a crucial component of credit assignment, the mixing network in value decomposition is essentially a state-conditioned feedforward neural network composed of multiple linear transformation layers and nonlinear activation functions. The network takes as input the individual Q-values of all agents, denoted as , and produces the joint Q-value as output. To enable the network to be state-dependent and adaptively parameterized, most value decomposition methods employ the hypernetwork mechanism [8]. Specifically, instead of being directly optimized, the weights and biases of each linear layer in the mixing network are generated by a hypernetwork that conditions on the global state . This design facilitates state-aware adaptation while ensuring compatibility with the centralized training and decentralized execution (CTDE) framework, making deployment straightforward. However, such neural-network-based credit assignment functions cannot effectively incorporate prior knowledge, which may result in slower convergence and reduced interpretability.
To address these issues, we propose Training-Free Credit Assignment Function (TFCAF), which is the core module of QLLM. This module is generated by LLMs through iterative error correction, based on a series of task-related prompts. As illustrated in Equation (2), the global Q-value function can theoretically be decomposed into a weighted sum of individual local Q-values with an additional bias term. Motivated by this decomposition paradigm, we incorporate a similar formulation in the design of TFCAF. Its mathematical expression is as follows:
The coefficient is a state-dependent weight associated with the global state , and is a bias term that also depends on the global state. Both terms constitute key components of the TFCAF, which are generated by LLMs. In the following, we provide a detailed explanation of how the TFCAF is generated and refined through error correction, as well as how it facilitates credit assignment in multi-agent systems.
4.2 Coder-Evaluator Framework
Due to the susceptibility of LLMs to hallucinations, the generated TFCAF must undergo continuous refinement. To address this challenge, we propose the coder-evaluator framework, which automates the generation of TFCAF through a modular and iterative process, ensuring both accuracy and robustness. This framework employs two LLMs with distinct yet complementary roles: the coder LLM , responsible for generating candidate TFCAFs, and the evaluator LLM , which critically evaluates these candidates and provides constructive feedback for improvement. The interaction between the two models is driven by two types of prompts: task prompts, which describe the reinforcement learning task in detail, and role prompts, which define the specific responsibilities and operational roles of each LLM within the framework. The role prompts are further categorized into coder prompts and evaluator prompts, corresponding to the Coder LLM and the Evaluator LLM, respectively. The iterative generation and refinement of TFCAF can be formally described through the following three clearly-defined steps.
Initial Generation.
Firstly, generates candidate credit assignment functions by processing both the role-specific and task-specific prompts:
where denote candidate TFCAFs, each including instructions for calculating state-dependent weights and biases. Given the inherent hallucination tendencies of LLMs, candidate functions might contain logical or dimensionality errors. To address this, each candidate TFCAF undergoes a validation step by testing it with state information and local Q-values from a representative environment scenario. If an error occurs (such as dimension mismatch or logical inconsistencies), the error message is fed back to , prompting regeneration of the faulty TFCAF:
This correction step repeats until all candidate TFCAFs pass basic execution checks without errors.
Error Detection and Correction.
Subsequently, evaluates these error-free candidate TFCAFs, selecting the most promising one, while also providing specific programming tips and insights for further optimization:
where . The selection and programming tips produced by are incorporated into a new prompt for , guiding it to generate another candidate functions:
where . Through repeated iterations of this coder-evaluator interaction, the candidate functions are progressively refined until convergence to an optimized final TFCAF . This final function explicitly includes a state-dependent weights module and a bias module .
Iterative Refinement.
Ultimately, is integrated into a predefined TFCAF template (see Appendix A), replacing the traditional mixing network and effectively performing credit assignment across multiple agents in MARL scenarios.
4.3 Training Procedure
With the coder-evaluator framework, we obtain a training-free credit assignment function. Like QMIX and VDN, QLLM is a value-based MARL method developed under the CTDE paradigm. Its objective is to learn a global action-value function that can be decomposed into individual utility functions , while avoiding the constraints imposed by conventional hypernetwork-based mixing networks. This design enables the LLM to model more complex and nonlinear interactions between local and global Q-values, allowing it to better adapt to the specific structural properties of the task.
As with most value decomposition approaches, each agent in QLLM employs a deep neural network to approximate its individual action-value function , where denotes the set of learnable parameters. These individual Q-values are then aggregated by the TFCAF, which is automatically generated by LLMs, to produce the global action-value function:
where is a parameterized function in which the weights are allowed to take arbitrary values, while the biases are also unconstrained. receives the agent-specific Q-values as input and utilizes the global state to generate the corresponding weights and biases through the weight module and the bias module , respectively. Notably, all parameters in the TFCAF are generated by LLMs and remain fixed during training. The global Q-value is then computed as shown in Equation (3). The training objective is to minimize the temporal-difference (TD) loss between the predicted total Q-value and the target Q-value:
where the target value is computed using Bellman equation:
with being the discount factor. denotes the parameters of the target agent network.
Compared to other neural network-based approaches for multi-agent credit assignment, QLLM involves fewer trainable parameters. This facilitates more precise credit allocation among agents in the early training stages, thereby accelerating the convergence of their decision-making policies. Meanwhile, the incorporation of TFCAF allows QLLM to effectively exploit prior knowledge while employing an interpretable credit assignment mechanism.
5 Experiments
To comprehensively evaluate the proposed method, the main experiments were designed with four main objectives: (1) to assess the performance of the QLLM algorithm in terms of average episode return and variance across multiple environments; (2) to verify its adaptability to high-dimensional global state spaces by increasing the dimensionality in selected tasks; (3) to demonstrate the compatibility of QLLM by integrating it into other value decomposition algorithms; and (4) to validate the effectiveness of the proposed coder-evaluator framework under various conditions.
5.1 Experimental Setups
We tested QLLM’s performance on four common MARL benchmarks and compared it with several well-known baseline credit assignment algorithms. The primary large language model used in our method is the DeepSeek-R1 [7] inference model. For each experiment, the reported return values represent the mean over five independent training runs, with 95% confidence intervals indicated by error bars. Unless otherwise specified, all environment parameters are set to their default values. Our experimental test environments include: (1) Cooperative Matrix Game111https://github.com/uoe-agents/matrix-games: This is a non-monotonic cooperative case where two cooperating agents must find an optimal Nash equilibrium [44]. (2) Multi-agent Particle Environments (MPE)222https://pettingzoo.farama.org/environments/mpe/: This environment consists of a series of cooperative tasks involving multiple particles, including navigation, predation, and communication. Agents must also avoid penalties caused by collisions. (3) Level-Based Foraging (LBF)333https://github.com/uoe-agents/lb-foraging: Multiple agents must collaborate to collect all the food items in the map. Successful collection occurs only when the sum of the levels of the participating agents is greater than or equal to the level of the food. (4) Robotic Warehouse (RWARE)444https://github.com/uoe-agents/robotic-warehouse: Robots continuously receive instructions to retrieve goods from specified locations and transport them to target locations. The detailed parameter settings, reward settings, and global state-space settings for each environment can be found in Appendix B.
We selected five commonly used baseline algorithms for comparison in this study:
Value Decomposition Algorithms: QMIX [28] is a value-based MARL algorithm that learns a joint Q-value function by factorizing individual agent value functions while maintaining a monotonic relationship between them. This allows agents to work independently, yet still guarantee optimal joint actions.
Policy Gradient Algorithms COMA [6] is a policy-gradient-based approach that uses counterfactual reasoning to address the credit assignment problem in multi-agent settings. It aims to improve the learning of individual agents by considering the effect of each agent’s action on the overall outcome. MADDPG [19] extends the DDPG algorithm to multi-agent settings. It utilizes centralized training with decentralized execution, where each agent learns a policy conditioned on the global state during training, but acts based on its local state during execution.
Reward Interpolation-Based Algorithms: RIIT [10] is a reward-based algorithm that improves multi-agent learning by interpolating rewards across agents. This helps agents better coordinate their actions to achieve a common objective.
Attention-Based Algorithms: Qatten [43] integrates attention mechanisms into Q-learning, allowing agents to dynamically adjust the focus on different aspects of the environment. This improves coordination and allows the agents to make more informed decisions.
5.2 The Superiority of QLLM
Average Return and Convergence Speed.
Figure 4 and Figure 5 shows the performance of various algorithms in certain scenarios. It is evident that the QLLM algorithm outperforms the baseline in both the dense-reward MPE environment and the sparse-reward LBF environment. Especially in the LBF environment, COMA, MADDPG, and RIIT almost fail to train the agents properly, which highlights the importance of accurate credit assignment in sparse-reward environments. In Listing 1, we present the credit assignment function generated by the large language model for the 2s-10x10-3p-3f map in the LBF environment. It is evident that this function takes into account key factors such as the positions of the agents and food, the effectiveness of the food, and the actions of the agents, accurately assigning contributions to each agent. Furthermore, it is easily extendable to similar scenarios.
Performance in Higher Dimensional State Spaces.
We also tested the performance of various algorithms in high-dimensional state spaces in certain scenarios. Taking the simple-spread task in MPE environment as an example, in the navigation scenario, as the number of agents increases, the global state space dimension becomes very large, and the difficulty of credit assignment grows exponentially. From the Figure 6, it can be seen that most algorithms perform poorly when faced with such a high-difficulty task, while our algorithm still maintains a high level of contribution assignment accuracy, which is attributed to the mechanism of using LLMs in the QLLM. The baseline algorithms rely entirely on neural networks for credit assignment, and as the state dimensions increase, determining the weight values becomes more difficult. However, our algorithm uses a set of well-defined nonlinear functions to assign contributions, which are unaffected by high-dimensionality, thus maintaining a high level of performance.
Compatibility Test.
Our algorithm is still based on value decomposition methods, so it is compatible with most MARL algorithms that use mixing networks to compute global Q-values. We replaced the mixing network in RIIT with the TFCAF generated by QLLM and tested the impact of using TFCAF on the performance of RIIT in three environment of MPE. The experimental results are shown in Table 1. Our results show that QLLM can be seamlessly incorporated into existing architectures, offering a flexible and scalable solution for a wide range of multi-agent settings.
Environment | RIIT-QLLM | RIIT |
pz-mpe-simple-spread-N6 | -169.7 3.4 | -201.1 2.7 |
pz-mpe-simple-adversary | 101.5 1.8 | 96.3 0.9 |
pz-mpe-simple-tag | 217.2 1.3 | 195.7 0.7 |
5.3 Ablation Studies
5.3.1 Coder-Evaluator Framework’s Candidate Responses
To study the impact of the number of candidate responses, a key hyperparameter in QLLM, Figure 7 shows the post-convergence average episode return of QLLM with varying on the tasks simple-tag and lbf:2s-10x10-3p-3f. The results indicate that as long as a sufficient number of candidate responses are provided, QLLM is insensitive to this hyperparameter. This finding supports our claim that the coder-evaluator framework helps reduce the randomness in LLM inference. Therefore, we set for all tasks in this work.
5.3.2 Evaluator LLM’s Effectiveness
Most RL algorithms enhanced by large language models use only one kind of LLM to do his job directly. Our coder-evaluator framework features two large language models, each with its specific role, working together to generate an accurate credit assignment function. Therefore, we need to explore in depth the effect of the evaluator LLM on experimental performance and investigate how it helps mitigate the issues of hallucination and randomness in large language models. We test the performance of QLLM across different reinforcement learning environments without using the evaluator LLM. Detailed experimental details can be found in Appendix B.
6 Conclusion
In this study, we propose QLLM, a method that leverages large language models to address the credit assignment problem in multi-agent reinforcement learning. Our algorithm employs a coder-evaluator framework to generate training-free, task-specific credit assignment functions. This frame significantly reduces hallucination and laziness in code generation by large language models, while also improving the accuracy of task understanding.
Extensive experiments conducted on several standard MARL benchmarks demonstrate that the proposed method consistently outperforms existing state-of-the-art baselines. Moreover, the proposed algorithm exhibits strong generalization capability and is compatible with various MARL algorithms that employ mixing networks, making it a promising and generalizable solution for complex multi-agent scenarios.
Limitations & Future Works: A current limitation of this work lies in the need to manually craft prompts for reinforcement learning tasks, which can become challenging due to the complexity of certain tasks. As LLM capabilities continue to advance, prompt design in QLLM is expected to become easier, requiring less task-specific information and further reducing manual effort. Future work will focus on expanding the applicability of the coder-evaluator framework and deploying the QLLM algorithm in real-world multi-robot systems.